2 men M1 and M2 and 3 women W1, W2 and W3 are trying for promotion in a firm to a single post which is vacant. Those of the same sex have equal probabilities of getting promotion but each man is twice as likely to get the promotion as any woman.

1.) Find the probability that woman gets the promotion

2.) If M2 and W2 are couple, find the probability that one of them gets the promotion


1
Expert's answer
2020-08-19T19:18:30-0400

Let p1 be the probability that a woman gets the promotion,

and let p2 be the probability that a man gets the promotion.

Then by conditon

3p1+2p2=1

and p2=2p1.

So, 3p1+4p1=1

7p1=1

p1=1/7, p2=2/7.

The probability that a woman gets the promotion is equal to 1/7.

The probability that one of W2 or M2 gets the promotion is equal to p1+p2:

p1+p2=1/7+2/7=3/7

So, the probability that one of W2 or M2 gets the promotion is equal to 3/7.